# PropaInsight: Toward Deeper Understanding of Propaganda in Terms of Techniques, Appeals, and Intent

### v1

@@ Abstract
Our experiments show that off-the-shelf LLMs struggle with propaganda analysis, but training with PropaGaze significantly improves performance. Fine-tuned Llama-7B-Chat achieves 203.4% higher text span IoU in technique identification and 66.2% higher BertScore in appeal analysis compared to 1-shot GPT-4-Turbo. Moreover, PropaGaze complements limited human-annotated data in data-sparse and cross-domain scenarios, showing its potential for comprehensive and generalizable propaganda analysis. [^1] PropaGaze and code will be released upon acceptance.
@@ 3.1 PTC-Gaze: Human-Annotated Dataset
PTC-Gaze builds on the existing PTC dataset Martino et al. (2020a), which includes human-written news articles annotated for propaganda techniques and spans. We reannotate this dataset by hiring human annotators to label appeals and intent independently. For appeals, annotators review propaganda-containing sentences along with their context and describe the feelings evoked. To reduce cognitive load, we provide GPT-4 generated candidate annotations for assistance. Annotators then evaluate whether the generated candidates accurately reflect their interpretations and reactions, and if not, they rewrite the descriptions based on the template in § 2 and Appendix C. For intent, annotators read the full article and infer its underlying intent in a single free-form sentence, and we leave the multi-intent scenarios for future work.As shown in Table 1, this annotated sub-dataset contains 79 articles, with an average of 12.77 propaganda techniques per article. Additional information, data examples, and analysis of the annotation quality are given in Appendix D.
@@ 4.1 Experimental Setup
In real-world scenarios, obtaining a large volume of well-annotated data for analyzing propaganda is challenging, as discussed in § 3. To simulate this limitation, we use both data-rich and data-sparse environments to evaluate the impact of larger, generalized synthetic datasets on model performance.
@@ 7 Conclusion and Future Work
    Although our dataset PropaGaze is novel and reliable, its size is relatively small due to the computational costs associated with GPT-4 and the high expense of human annotation. Consequently, we are uncertain about the dataset’s ability to generalize across a broad range of domains when models are fine-tuned exclusively on it.
@@ Acknowledgement
This research was done with funding from the Defense Advanced Research Projects Agency (DARPA) under Contract No. HR001120C0123. The views, opinions, and/or findings expressed are those of the authors and should not be interpreted as representing the official views or policies of the Department of Defense or the U.S. Government. We also thank Kitware.Inc. and Rapidata.Inc. for their help in the data annotation process.
@@ Appendix B Data Generation Prompt Templates
In this task, the input will be a news article, then some sentence in the article will be provided and you need to identify how the specific sentence raise appeals among the readers, the propaganda tactics used in these sentences will also be proivded as a hint. Also remeber that your response should be aware of the main goal of the whole article. For each sentence, you only need to output a sentence describing the feelings in one of the following two templates:

### v2

@@ Abstract
Our experiments show that off-the-shelf LLMs struggle with propaganda analysis, but training with PropaGaze significantly improves performance. Fine-tuned Llama-7B-Chat achieves 203.4% higher text span IoU in technique identification and 66.2% higher BertScore in appeal analysis compared to 1-shot GPT-4-Turbo. Moreover, PropaGaze complements limited human-annotated data in data-sparse and cross-domain scenarios, showing its potential for comprehensive and generalizable propaganda analysis. [^1] PropaGaze and code are available at this GitHub repository: https://github.com/Lumos-Jiateng/PropaInsight.
@@ 3.1 PTC-Gaze: Human-Annotated Dataset
PTC-Gaze builds on the existing PTC dataset Martino et al. (2020a), which includes human-written news articles annotated for propaganda techniques and spans. We reannotate this dataset by hiring human annotators to label appeals and intent independently. For appeals, annotators review propaganda-containing sentences along with their context and describe the feelings evoked. To reduce cognitive load, we provide GPT-4 generated candidate annotations for assistance. Annotators then evaluate whether the generated candidates accurately reflect their interpretations and reactions, and if not, they rewrite the descriptions based on the template in § 2 and Appendix C. For intent, annotators read the full article and infer its underlying intent in a single free-form sentence, and we leave the multi-intent scenarios for future work.We used Label Studio to design the annotation interface, which is later shown in Appendix D. Two professional annotators from Kitware Inc. are in charge of the annotation task. We only have one annotator for each annotation task so no agreement rate is computed. As shown inTable 1, this annotated sub-dataset contains 79 articles, with an average of 12.77 propaganda techniques per article. Additional information, data examples, and analysis of the annotation quality are given in Appendix D.
@@ 4.1 Experimental Setup
In real-world scenarios, obtaining a large volume of well-annotated data for analyzing propaganda is challenging, as discussed in § 3. 6tt
@@ 7 Conclusion and Future Work
    Although our dataset, PropaGaze, is novel and reliable, its size is relatively small due to the computational costs associated with GPT-4 and the high expense of human annotation. Consequently, we are uncertain about the dataset’s ability to generalize across a broad range of domains when models are fine-tuned exclusively on it.
@@
  5. 5.
    Our research is currently limited to English, which may restrict the generalizability of our findings to other languages. Future work will focus on extending the approach to cross-lingual settings to address this limitation.
  6. 6.
    We acknowledge that while our framework attempts to model intrinsic aspects of propaganda, the experimental setup simplifies the characterization of propaganda intentions. Specifically, the process of generating and validating intentions relies on annotator feedback rather than leveraging domain-specific intent modeling, which may limit the system’s ability to fully capture nuanced propaganda strategies.
@@
  6. 6.
    We plan to incorporate intent mining approaches that include domain-specific characterizations of propaganda intentions. By integrating techniques such as hierarchical intent modeling, pretrained language models fine-tuned for intent detection, and emotion-informed analysis, we aim to strengthen our framework’s ability to detect and classify propaganda intentions with greater specificity and robustness.
@@ Acknowledgement
This research was done with funding from the Defense Advanced Research Projects Agency (DARPA) under Contracts No. HR001120C0123 and HR0011-24-3-0325. The views, opinions, and/or findings expressed are those of the authors and should not be interpreted as representing the official views or policies of the Department of Defense or the U.S. Government. We also thank Kitware.Inc. and Rapidata.Inc. for their help in the data annotation process.
@@ Appendix B Data Generation Prompt Templates
In this task, the input will be a news article, then some sentence in the article will be provided and you need to identify how the specific sentence raises appeals among the readers, the propaganda tactics used in these sentences will also be proivded as a hint. Also remeber that your response should be aware of the main goal of the whole article. For each sentence, you only need to output a sentence describing the feelings in one of the following two templates: